OpenStudy (anonymous):

The length of a rectangle is 4 inches more than its width, and its perimeter is 34 inches. Find the length and width of the rectangle. Find the width of the rectangle.

OpenStudy (anonymous):

width=6

OpenStudy (anonymous):

Width=6.5.

OpenStudy (anonymous):

The length is x+4, and the width is x, because the length, or L, is 4 more than the width, or W. This is what the equation would look like at first: x+x+x+4+x+4=34. Simplify the equation: add all of the variables on the left side. 4x+8=34. Subtract 8 from each side: 4x=26. Divide 4 from each side: x=6.5. That's the width.

OpenStudy (anonymous):

The length is x+4, or 6.5 +4. which is 10.5.

OpenStudy (anonymous):

Here is how to prove it: all of the lengths of the 4 sides should make up 34. Let's see. 6.5+6.5=13. 10.5+10.5=21. 21+13=34. The answer is correct.